# Some remarks on Gorenstein projective precovers

### v2

@@ Abstract.
In this paper, we prove that for a $n$-perfect ring $R$ various classes of relative Gorenstein projective $R$-modules are special precovering, among them including the Gorenstein projectives and Ding projectives. We got a family of hereditary and complete cotorsion pairs and hereditary Hovey triples that comes from relative Gorenstein projective $R$-modules. We also introduce a new class of Gorenstein projective $R$-modules relative to the Auslander class $\mathcal{A}_{C}(R)$ of a semidualizing $(R,S)$-bimodule ${}_{R}C_{S}$.
@@ 1. Introduction
A central question in homological theory of modules is to understand the size of the homological dimensions which are associated with rings and modules. In particular the famous second finitistic conjecture is established from the question of how large can the projective dimension of modules over a ring be?. To address this concern consider a ring $R$ and $\mathcal{P}(R)^{<\infty}_{fin}$ the class of $R$-modules finitely generated of finite projective dimension, the Second Finitistic Dimension Conjeture establish that $fpd(R):=\sup\{\mathrm{pd}_{R}(M)|M\in\mathcal{P}(R)^{<\infty}_{fin}\}$ is finte.
@@
It is well-known that the understanding of the class $\mathcal{P}(R)^{<\infty}_{fin}$ and $\mathcal{P}(R)^{<\infty}$ provides insight to finitistic dimension conjectures as was proved by Auslander and Reiten in [1] who proved that contravariant finiteness of the class [^1] In this paper we use precovering as synonym of contravariantly finite. $\mathcal{P}(R)^{<\infty}_{fin}$, referred to as the Auslander–Reiten condition from now on, is a sufficient condition for validity of the second finitistic dimension conjecture over an Artin algebra. More recently it is proved by P. Moradifar and J. Šaroch [21] that contravariant finiteness of the class $\mathcal{GP}^{<\infty}_{fin}$ (finitely generated modules of finite Gorenstein projective dimension) implies validity of the second finitistic dimension conjecture over left artinian rings. Furthermore is proved that that contravariant finiteness of the class $\mathcal{GP}(R)^{<\infty}_{fin}$ implies contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$, over rings where $\mathcal{GP}(R)_{fin}$ (the class of finitely generated Gorenstein projective modules) is contravariantly finite and the converse holds for Artin algebras. In a more general context S. Estrada, A. Iacob and K. Yeomans have proven that the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$ provided that the ring $R$ be GF-closed, right coherent and left $n$-perfect [13, Theorem 2]. And more recently, a relationship has been found between the existence of Gorenstein projective precovers and finitely presented $R$-modules with the Second Finitistic Dimension Conjeture [11, Theorem 5]. Thus, in view of the importance contravariant finiteness of the class $\mathcal{GP}(R)$, we study when such class and their generalizations are special precovering on $\mathrm{Mod}(R)$. In this paper we prove in Proposition 3.2 and Corollary 4.8 that when $R$ is left $n$-perfect and $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$ the classes of Gorenstein projectives $\mathcal{GP}(R)$ and Ding projectives $\mathcal{DP}(R)$ are special precovering. Also in Proposition 3.4 we prove that the class of projectively coresolved Gorenstein flat $R$-modules $\mathcal{P}\mathcal{GF}(R)$ is special precovering provided that $R$ is left $n$-perfect. We then turn our attention to the classes of Gorenstein projective $R$-modules that come from a duality pair $(\mathcal{L},\mathcal{A})$ and we prove in Proposition 4.9 that for a complete duality pair over a $n$-perfect ring $R$, the class of Gorenstein projectives $\mathcal{GP}_{\mathcal{L}}(R)$ relative to this pair $(\mathcal{L},\mathcal{A})$ is special precovering. As an application, we recover in Proposition 4.11 the main result of [13]. In Theorem 4.14 we get families of cotorsion pairs and in Corollary 4.15 families of Hovey triples. At the end in Proposition 4.16 we declare an application of the developed theory on a new class of Gorenstein projective $R$-modules.
@@ 2. Preliminaries
Concerning functors defined on modules, $\mathrm{Ext}^{i}_{R}(-,-)$ denotes the right $i$-th derived functor of $\mathrm{Hom}_{R}(-,-)$. If $M\in\mathrm{Mod}(R^{\rm op})$ and $N\in\mathrm{Mod}(R)$, $M\otimes_{R}N$ denotes the tensor product of $M$ and $N$. Recall the construction of this tensor products defines a bifunctor $-\otimes_{R}-\colon\mathrm{Mod}(R^{\rm op})\times\mathrm{Mod}(R)\longrightarrow\mathrm{Mod}(\mathbb{Z})$, where $\mathrm{Mod}(\mathbb{Z})$ is the category of abelian groups.
@@
$\mathcal{X}^{\perp_{i}}$ :=$:=\{M\in\mathrm{Mod}(R){\rm\ :\ }\mathrm{Ext}^{i}_{R}(\mathcal{X},M)=0\}$ and $\mathcal{X}^{\perp}$ $:=\bigcap_{i\geq 1}\mathcal{X}^{\perp_{i}}.$
@@
$\mathrm{resdim}_{\mathcal{X}}(M):=\min\{m\in\mathbb{Z}_{\geq 0}\ \mbox{ : }\ \text{there exists an $\mathcal{X}$-resolution of $M$ of length $m$}\}.$
@@
$\mathcal{X}^{\wedge}_{n}$ :=$:=\{M\in\mathrm{Mod}(R)\text{ : }\mathrm{resdim}_{\mathcal{X}}(M)\leq n\}$ and $\mathcal{X}^{\wedge}$ $:=\bigcup_{n\geq 0}\mathcal{X}^{\wedge}_{n}.$
@@ 3. Gorenstein projective precovers
From Šaroch and Št’ovíchěk’s [23] we know now that every ring is GF-closed [^2] This means that $\mathcal{GF}(R)$ is closed by extensions for all ring $R$.. Thus, is enough that the ring $R$ will be right coherent and left $n$-perfect [^3] This means that for the ring $R$ and $n\geq 0$ the containment $\mathcal{P}(R)\subseteq\mathcal{F}(R)^{\wedge}_{n}$ is given. to obtain $\mathcal{GP}(R)$ special precovering. In what follows we will see that such result remains true under other conditions less known and that the above result can be obtained for classes of relative Gorenstein projective $R$-modules.
@@
Consider $R$ a left $n$-perfect ring. Then, every Gorenstein flat $R$-module has a $\mathcal{GP}(R)$-resolution at most $n$. That is, the containment $\mathcal{GF}(R)\subseteq\mathcal{GP}(R)^{\wedge}_{n}$ is given.
@@
Indeed, from [12, Proposition 3.1 (1)] the condition $\mathrm{Ext}_{R}^{i}(G,F)=0$ for all $G\in\mathcal{GF}(R)$ and $F\in\mathcal{F}\cap\mathcal{F}^{\bot}$ is true without conditions over the ring. Thus by using [13, Proposition 2] we conclude that $\mathcal{GF}(R)\subseteq\mathcal{GP}(R)^{\wedge}_{n}$. ∎
We have here our first main result.
Proposition 3.2.
Let $R$ be a ring left $n$-perfect and such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$. Then the class $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$.
Proof.
We will verify the conditions of [13, Theorem 1]. From [27, Corollary 2.6] we have the containment $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$, since $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$. While the containment $\mathcal{GF}(R)\subseteq\mathcal{GP}(R)^{\wedge}_{n}$ it follows from Proposition 3.1. ∎
This situation also arises for the class of projectively coresolved Gorenstein flat $R$-modules $\mathcal{P}\mathcal{GF}(R)$ introduced by Šaroch and Št’ovíchěk [23]. Such class $\mathcal{P}\mathcal{GF}(R)$ consist of cycles of exact complexes of left projective $R$-modules which remains exact after applying the funtor $I\otimes_{R}-$, for all $I\in\mathcal{I}(R^{\mathrm{op}})$. To see how this happens, we need a useful tool, which is more extensive than the one presented in [9, Proposition 3.1]. The notation $P\mathrm{Gfd}(M)$ means $\mathrm{resdim}_{\mathcal{P}\mathcal{GF}}(M)$, called the projectively coresolved Gorenstein flat dimension of $M$.
Proposition 3.3.
Let $R$ be a ring and consider $\mathcal{P}\mathcal{GF}(R)$ the class of projectively coresolved Gorenstein flat $R$-modules. Then for all $M\in\mathrm{Mod}(R)$ with $P\mathrm{Gfd}(M)=n<\infty$, there are exact sequences in $\mathrm{Mod}(R)$
$0\to K\to X\stackrel{{\scriptstyle\varphi}}{{\to}}M\to 0,$
with $\mathrm{pd}(K)=n-1$ and $X\in\mathcal{P}\mathcal{GF}(R)$, where $\varphi:X\to M$ is an $\mathcal{P}\mathcal{GF}(R)$-precover and $K\in\mathcal{P}\mathcal{GF}(R)^{\perp}$.
And
$0\to M\stackrel{{\scriptstyle\varphi^{\prime}}}{{\to}}H\to X^{\prime}\to 0,$
with $\mathrm{pd}(H)=n$ and $X^{\prime}\in\mathcal{P}\mathcal{GF}(R).$ Where $\varphi^{\prime}:M\to H$ is an $\mathcal{P}(R)^{\wedge}$-preenvelope and $X^{\prime}\in{}^{\perp}(\mathcal{P}(R)^{\wedge})$.
Proof.
Since $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{GP}(R)$ [23, Theorem 4.4] and the pair $(\mathcal{GP}(R),\mathcal{P}(R))$ is a strong left Frobenius pair then the pair $(\mathcal{P}\mathcal{GF}(R),\mathcal{P}(R))$ forms also a strong left Frobenius pair, since $\mathcal{P}\mathcal{GF}(R)$ is projective resolving closed under direct sums and direct summands. Thus, the result it follows from apply [6, Theorem 2.8]. ∎
We have here our second main result.
Proposition 3.4.
Let $R$ be a left $n$-perfect ring, then the class of projectively coresolved Gorenstein flat $R$-modules $\mathcal{P}\mathcal{GF}(R)$ is special precovering on $\mathrm{Mod}(R)$.
Proof.
Actually we know that $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{GF}(R)$, without conditions on $R$. Now, from [9, Proposition 3.9], we have that $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}_{n}$. We will use this in the proof, as follows. Let $X\in\mathrm{Mod}(R)$, by [14, Theorem 3.4] we know that $(\mathcal{GF}(R),\mathcal{GF}(R)^{\bot})$ is a hereditary and complete cotorsion pair. Thus, there is an exact sequence $0\to Y\to N\to X\to 0$ with $N\in\mathcal{GF}(R)$ and $Y\in\mathcal{GF}(R)^{\bot}$. From the containment $\mathcal{P}\mathcal{GF}(R)\subseteq\mathcal{GF}(R)$, we get $\mathcal{GF}(R)^{\bot}\subseteq\mathcal{P}\mathcal{GF}(R)^{\bot}$. Since $\mathcal{GF}(R)\subseteq\mathcal{P}\mathcal{GP}(R)^{\wedge}_{n}$. By Proposition 3.3 for $N\in\mathcal{GF}(R)$ there is an exact sequence $0\to W\to T\to N\to 0$ with $T\in\mathcal{P}\mathcal{GF}(R)$ and $W\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. We can construct the following p.b digram
[…]
to obtain the exact sequence $0\to W\to A\to Y\to 0$, with $W,Y\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. This implies that $A\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. Therefore we obtain the sequence $0\to A\to T\to X\to 0$, with $T\in\mathcal{P}\mathcal{GF}(R)$ and $A\in\mathcal{P}\mathcal{GF}(R)^{\bot}$. Therefore $T\to X$ is a special $\mathcal{P}\mathcal{GF}(R)$-precover. ∎
4. Generalized Gorenstein flat modules
@@
In what follows, we shall consider classes $\mathcal{X}\subseteq\mathrm{Mod}(R)$ and $\mathcal{Y}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. The following definition of Gorenstein flat $R$-modules relative to a pair $(\mathcal{X},\mathcal{Y})$ comes from [28, Definition 2.1].
@@
An $R$-module $M$ is Gorenstein $(\mathcal{X},\mathcal{Y})$-flat if there exists an exact and $(\mathcal{Y}\otimes_{R}-)$-acyclic complex $X_{\bullet}\in\mathrm{Ch}(\mathcal{X})$ such that $M\simeq Z_{0}(X_{\bullet})$. By $(\mathcal{Y}\otimes_{R}-)$-acyclic we mean that $Y\otimes_{R}X_{\bullet}$ is an exact complex of abelian groups for every $Y\in\mathcal{Y}$. The class of Gorenstein $(\mathcal{X,Y})$-flat $R$-modules will be denoted by $\mathcal{GF}_{(\mathcal{X,Y})}$.
There are several types of classes Gorenstein $(\mathcal{X},\mathcal{Y})$-flat that have been presented in recent literature. In [23, §4] is presented by J. Šaroch and J. Št’ovíček, the class of projectively coresolved Gorenstein flat $R$-modules, here denoted $\mathcal{P}\mathcal{GF}(R)$, later generalized and extensively studied by S. Estrada, A. Iacob and M. A. Pérez in [12, Definition 2.6]. This class is called projectively coresolved Gorenstein $\mathcal{B}$-flat and here denoted $\mathcal{P}\mathcal{GF}_{\mathcal{B}}$. In such paper also is studied the class $\mathcal{GF}_{\mathcal{B}}$ (in the notation of Definition 4.1 is $\mathcal{GF}_{(\mathcal{F},\mathcal{B})}$ where the subscript $\mathcal{F}$ denotes $\mathcal{F}(R)$) with $\mathcal{B}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$ is sometimes a semi-definable class [12, Definition 2.7].
In a different environment is defined by J. Gillespie in [16] for a complete duality pair $(\mathcal{L},\mathcal{A})$ the class of Gorenstein $(\mathcal{L},\mathcal{A})$-flat $R$-modules, here denoted by $\mathcal{GF}_{\mathcal{A}}$, and is studied [16, §5] some of their model structures. In such paper also is defined the class of Gorenstein $(\mathcal{L},\mathcal{A})$-projective $R$-modules here denoted by $\mathcal{GP}_{\mathcal{L}}$ and the class of Gorenstein $(\mathcal{L},\mathcal{A})$-injective $R$-modules here denoted by $\mathcal{GI}_{\mathcal{A}}$. We see that when $(\mathcal{L},\mathcal{A})$ is a duality pair the classes $\mathcal{GP}_{\mathcal{L}}$, $\mathcal{GI}_{\mathcal{A},}$, $\mathcal{GF}_{\mathcal{A}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ have an interesting interaction between them (see [4], [12], [16], [17], [26]). In fact the classes $\mathcal{GP}_{\mathcal{L}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ agree sometimes. This is true for a symmetric duality pair $(\mathcal{L},\mathcal{A})$, from [7, Theorem A.6.] [^4] Note that for D. Bravo et. al, the notion of duality pair differs from the one presented here. since the class $\mathcal{L}$ is closed by pure quotients [18, Theorem 3.1]. This fact had been mentioned previously by J. Gillespie and A. Iacob in [17, Corollary 14]. We declare this result as follows.
@@
When for a duality pair $(\mathcal{L},\mathcal{A})$ the class $\mathcal{A}$ is semi-definable then $\mathcal{GP}_{\mathcal{L}}$ is special precovering in the whole category $\mathrm{Mod}(R)$, since [12, Theorem 2.13] the class $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ is the left part of a complete and hereditary cotorsion pair. A duality pair with such property is the definable pair mentioned in [17, Example 12] and denoted $(<\mathcal{F}(R)>,<\mathcal{I}(R^{\mathrm{op}})>)$. That is, the class of Gorenstein $(\mathcal{P}(R),<\mathcal{F}(R)>)$-projective $R$-modules is special precovering in $\mathrm{Mod}(R)$. We state this as follows.
@@
Consider the definable duality pair $(<\mathcal{F}(R)>,<\mathcal{I}(R^{\mathrm{op}})>)$ in $\mathrm{Mod}(R)$, then the class $\mathcal{GP}_{<\mathcal{F}(R)>}$ is the left part of a complete and hereditary cotorsion pair.
For each $n\geq 2$ there is a duality pair $(\mathcal{FP}_{n}\mbox{-}Flat(R),\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}}))$, which consists of the $FP_{n}$-flat left R-modules and $FP_{n}$-injective right $R$-modules [8, Theorems 5.5 & 5.6]. The right part $\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}})$ is definable [12, Example 2.21. (3)]. The classes of Gorenstein $R$-modules that comes from these duality pairs have been studied in [20] by A. Iacob, proving in a manner similar to that used in Proposition 4.3 that $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ is the left part of a complete and hereditary cotorsion pair, and thus is a class special precovering. From here, we have the following result.

### v3

@@ Abstract.
The existence of the Gorenstein projective precovers over arbitrary rings is an open question. In this paper, we make use of three different techniques addressing intrinsic and homological properties of several classes of relative Gorenstein projective $R$-modules, among them including the Gorenstein projectives and Ding projectives, with the purpose of giving some situations where Gorenstein projective precovers exists. Within the development of such techniques we obtaint a family of hereditary and complete cotorsion pairs and hereditary Hovey triples that comes from relative Gorenstein projective $R$-modules. We also study a class of Gorenstein projective $R$-modules relative to the Auslander class $\mathcal{A}_{C}(R)$ of a semidualizing $(R,S)$-bimodule $_{R}C_{S}$, where we make use of a property of “reduction”.
@@ 1. Introduction
For $R$ an associative ring with identity, the Gorenstein projective, injective and flat $R$-modules where introduced in [14], since then the Gorenstein homological algebra has been developing intensively as a relative version of homological algebra that replaces the classical projective, injective and flat modules and resolutions with the Gorenstein versions. While in the classical homological algebra is know the existence of projective resolutions (resp. injective and flat resolutions) for any $R$-module without restrictions over $R$, the situation is different for the Gorenstein version. The question: What is the most general class of rings over which all modules have Gorenstein projective (injective) resolutions? still open. A comprehensive response has been given by S. Estrada, A. Iacob and K. Yeomans, who have proved that the class of Gorenstein projective $R$-modules $\mathcal{GP}(R)$ is special precovering on $\mathrm{Mod}(R)$ provided that the ring $R$ be right coherent and left $n$-perfect [19, Theorem 2]. More recently, a relationship has been found between the existence of Gorenstein projective precovers and finitely presented $R$-modules with the Second Finitistic Dimension Conjeture [17, Theorem 5]. Furthermore, a closer relationship has been presented by P. Moradifar and J. Šaroch in [29] where is proved that contravariant finiteness of the class [^1] In this paper we use precovering as synonym of contravariantly finite. $\mathcal{GP}(R)^{<\infty}_{fin}$ (finitely generated $R$-modules of finite Gorenstein projective dimension) implies validity of the Second Finitistic Dimension Conjeture over left artinian rings. Furthermore is proved that that contravariant finiteness of the class $\mathcal{GP}(R)^{<\infty}_{fin}$ implies contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$, over rings where $\mathcal{GP}(R)_{fin}$ (the class of finitely generated Gorenstein projective $R$-modules) is contravariantly finite and the converse holds for Artin algebras. Such relation is important, since was proved by Auslander and Reiten in [1] that contravariant finiteness of the class $\mathcal{P}(R)^{<\infty}_{fin}$, referred to as the Auslander–Reiten condition, is a sufficient condition for validity of the Second Finitistic Dimension Conjeture over an Artin algebra.
In view of the importance contravariant finiteness of the class $\mathcal{GP}(R)$, we study when such class is precovering on $\mathrm{Mod}(R)$. Although we know by M. Cortés-Izurdiaga, J. Šaroch [10] that the pair $(\mathcal{GP}(R),\mathcal{GP}(R)^{\bot})$ is always an hereditary cotorsion pair, a condition to be complete is that all projective modules are $\lambda$-pure-injective for some infinite regular cardinal $\lambda$ (in particular, if R is right $\Sigma$-pure-injective). In this paper we make use of three different ways to obtain $\mathcal{GP}(R)$ precovering. First, we make use of the intrinsic properties of $\mathcal{GP}(R)$ to see when the Holm’s question [13, Remark 4.5 (4)] is fulfilled. For the second, we make use of the different generalizations of the class $\mathcal{GP}(R)$ analyzing when these classes match. For the third, we make use of the technique of S. Estrada and A. Iacob [17], by finding a suitable complete hereditary cotorsion pair.
@@ 2. Preliminaries
Concerning functors defined on modules, $\mathrm{Ext}^{i}_{R}(-,-)$ denotes the right $i$-th derived functor of $\mathrm{Hom}_{R}(-,-)$. If $M\in\mathrm{Mod}(R^{\rm op})$ and $N\in\mathrm{Mod}(R)$, $M\otimes_{R}N$ denotes the tensor product of $M$ and $N$. Recall the construction of this tensor products defines a bifunctor $\otimes_{R}-\colon\mathrm{Mod}(R^{\rm op})\times\mathrm{Mod}(R)\longrightarrow\mathrm{Mod}(\mathbb{Z})$, where $\mathrm{Mod}(\mathbb{Z})$ is the category of abelian groups.
@@
$\mathcal{X}^{\perp_{i}}$ $:=\{M\in\mathrm{Mod}(R){\rm\ :\ }\mathrm{Ext}^{i}_{R}(\mathcal{X},M)=0\}$ and $\mathcal{X}^{\perp}$ $:=\bigcap_{i\geq 1}\mathcal{X}^{\perp_{i}}.$
@@
[…]
@@
$\mathcal{X}^{\wedge}_{n}$ $:=\{M\in\mathrm{Mod}(R)\text{ : }\mathrm{resdim}_{\mathcal{X}}(M)\leq n\}$ and $\mathcal{X}^{\wedge}$ $:=\bigcup_{n\geq 0}\mathcal{X}^{\wedge}_{n}.$
@@ 3. Gorenstein projective precovers
From Šaroch and Št’ovíchěk’s [31, Remark in p.24-25] we know now that the condition $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$ is true if and only if $\mathcal{GP}(R)=\mathcal{P}\mathcal{GF}(R)$, which, in a way, answers Holm’s question [13, Remark 4.5 (4)]. Where the class here denoted $\mathcal{P}\mathcal{GF}(R)$ is the presented in [31, §4] called the class of projectively coresolved Gorenstein flat $R$-modules which consist of cycles of exact complexes of left projective $R$-modules that remains exact after applying the functor $(I\otimes_{R}-)$ for all $I\in\mathcal{I}(R^{\mathrm{op}})$. Now, from [31, Theorem 4.9] the class $\mathcal{P}\mathcal{GF}(R)$ is always a special precovering class. We declare this situation as follows.
Remark 3.1.
Let $R$ be a ring such that $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)$, then the class $\mathcal{GP}(R)$ is special precovering.
From the previous result we can see that the conditions asked in [19, Theorem 1] there are more than needed. In what follows we address a variety of conditions that make $\mathcal{GP}(R)$ a class special precovering in $\mathrm{Mod}(R)$, for do this we use other kinds of relative projective Gorenstein modules and other particular notions that have recently appeared in the literature..
The statements (i)-(iv) in the following Proposition are basically equivalent to Holm’s question. Their proof it follows from [35, Theorem 2.4, Corollaries 2.5, 2,6] and Remark 3.1. It may be noted that there are other conditions involving the copure dimensions defined by Enochs and Jenda [14], which we do not deal with here.
@@
Given a ring $R$ the class $\mathcal{GP}(R)$ is special precovering in each one of the following situations:
  • (i)
    $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R^{\mathrm{op}})^{\wedge}$,
  • (ii)
    $\mathcal{GP}(R)\subseteq\mathcal{P}\mathcal{GF}(R)^{\wedge}$,
  • (iii)
    $\mathcal{GP}(R)\subseteq\mathcal{GF}(R)^{\wedge}$,
  • (iv)
    There exist an integer $n\geq 0$ such that $\mathrm{Tor}^{R}_{n}(I,M)=0$ for any $I\in\mathcal{I}(R^{\mathrm{op}})$ and $M\in\mathcal{GP}(R)$,
  • (v)
    If $R$ is either a right coherent ring such that $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$ or a ring such that $\mathcal{I}(R^{\mathrm{op}})\subseteq\mathcal{F}(R)^{\wedge}$.
Note that a situation where $R$ is close to being right coherent and left $n$-perfect  [^2] This means that for the ring $R$ and $n\geq 0$ the containment $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}_{n}$ is given. [19, Theorem 2] arises in Proposition 3.2 (v). In the following result we give a proof with another arguments where also $\mathcal{GP}(R)$ is special precovering. The purpose of giving the details is that the arguments can be adapted to other situations, where some kind of relative Gorenstein projective $R$-modules match with $\mathcal{GP}(R)$.
Proposition 3.3.
Let $R$ be a right coherent ring, then the class of Ding projective $R$-modules $\mathcal{DP}(R)$ is special precovering on $\mathrm{Mod}(R)$. Furthermore, if $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$ then $\mathcal{GP}(R)$ is also special precovering on $\mathrm{Mod}(R)$.
@@
Indeed, since $R$ is right coherent, the pair $(\mathcal{F}(R),\mathcal{I}(R^{\mathrm{op}}))$ is a symmetric duality pair, thus from [7, Theorem A.6.] [^3] Note that for D. Bravo et. al. the notion of duality pair differs from the one presented here. we get that $\mathcal{P}\mathcal{GF}(R)=\mathcal{DP}(R)$. Now from [31, Theorem 4.9] the class $\mathcal{P}\mathcal{GF}(R)$ is always a special precovering class, thus $\mathcal{DP}(R)$ is special precovering on $\mathrm{Mod}(R)$. Finally, if $\mathcal{F}(R)\subseteq\mathcal{P}(R)^{\wedge}$, we get from [4, Proposition 6.7] that $\mathcal{DP}(R)$ and $\mathcal{GP}(R)$ coincide, therefore $\mathcal{GP}(R)$ is special precovering. ∎
As is known, the class of Ding projectives appears as a generalization of Gorenstein projectives, while the class $\mathcal{P}\mathcal{GF}(R)$ comes to complement $\mathcal{GP}(R)$ since $\mathcal{P}\mathcal{GF}(R)$ is a subclass of $\mathcal{GP}(R)$ which consists of Gorenstein flat $R$-modules, which makes it more versatile, for example to show that every ring is GF-closed [31]. Thus, the generalizations and variants of $\mathcal{GP}(R)$ and $\mathcal{GF}(R)$ provide new information about them [28]. this is partly the reason why we address generalizations of $\mathcal{GP}(R)$ and $\mathcal{GF}(R)$ in order to provide a better understanding of them.
3.1. Precovers from generalized Gorenstein $R$-modules
@@
In what follows, we shall consider classes $\mathcal{X}\subseteq\mathrm{Mod}(R)$ and $\mathcal{Y}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$. The following definition of Gorenstein flat $R$-modules relative to a pair $(\mathcal{X},\mathcal{Y})$ comes from [36, Definition 2.1].
@@
An $R$-module $M$ is Gorenstein $(\mathcal{X},\mathcal{Y})$-flat if there exists an exact and $(\mathcal{Y}\otimes_{R}-)$-acyclic complex $X_{\bullet}\in\mathrm{Ch}(\mathcal{X})$ such that $M\cong Z_{0}(X_{\bullet})$. By $(\mathcal{Y}\otimes_{R}-)$-acyclic we mean that $Y\otimes_{R}X_{\bullet}$ is an exact complex of abelian groups for every $Y\in\mathcal{Y}$. The class of Gorenstein $(\mathcal{X,Y})$-flat $R$-modules will be denoted by $\mathcal{GF}_{(\mathcal{X,Y})}$.
There are several types of classes Gorenstein $(\mathcal{X},\mathcal{Y})$-flat that have been presented in recent literature. The class $\mathcal{P}\mathcal{GF}$ presentes above have been generalized and extensively studied by S. Estrada, A. Iacob and M. A. Pérez in [18, Definition 2.6]. This class is called projectively coresolved Gorenstein $\mathcal{B}$-flat and here denoted $\mathcal{P}\mathcal{GF}_{\mathcal{B}}$. In such paper also is studied the class $\mathcal{GF}_{\mathcal{B}}$ (in the notation of Definition 3.4 is $\mathcal{GF}_{(\mathcal{F},\mathcal{B})}$ where the subscript $\mathcal{F}$ denotes $\mathcal{F}(R)$) where $\mathcal{B}\subseteq\mathrm{Mod}(R^{\mathrm{op}})$ is sometimes a semi-definable class [18, Definition 2.7].
In a near environment, is defined by J. Gillespie in [22] for a complete duality pair $(\mathcal{L},\mathcal{A})$ the class of Gorenstein $(\mathcal{L},\mathcal{A})$-flat $R$-modules, here denoted by $\mathcal{GF}_{\mathcal{A}}$, and is studied [22, §5] some of their dimensions and model structures. In such paper also is defined the class of Gorenstein $(\mathcal{L},\mathcal{A})$-projective $R$-modules here denoted by $\mathcal{GP}_{\mathcal{L}}$ and the class of Gorenstein $(\mathcal{L},\mathcal{A})$-injective $R$-modules here denoted by $\mathcal{GI}_{\mathcal{A}}$. We see that when $(\mathcal{L},\mathcal{A})$ is a duality pair the classes $\mathcal{GP}_{\mathcal{L}}$, $\mathcal{GI}_{\mathcal{A},}$, $\mathcal{GF}_{\mathcal{A}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ have an interesting interaction between them (see [4], [18], [22], [23], [34]). In fact ,the classes $\mathcal{GP}_{\mathcal{L}}$ and $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ agree sometimes. This is true for a symmetric duality pair $(\mathcal{L},\mathcal{A})$, from [7, Theorem A.6.] since the class $\mathcal{L}$ is closed by pure quotients [24, Theorem 3.1]. This fact had been mentioned previously by J. Gillespie and A. Iacob in [23, Corollary 14]. We declare this result as follows.
@@
When for a symmetric duality pair $(\mathcal{L},\mathcal{A})$ the class $\mathcal{A}$ is semi-definable then $\mathcal{GP}_{\mathcal{L}}$ is special precovering in the whole category $\mathrm{Mod}(R)$, since [18, Theorem 2.13] the class $\mathcal{P}\mathcal{GF}_{\mathcal{A}}$ is the left part of a complete and hereditary cotorsion pair. A duality pair with such property is the definable pair mentioned in [23, Example 12], denoted $(\langle\mathcal{F}(R)\rangle,\langle\mathcal{I}(R^{\mathrm{op}})\rangle)$. That is, the class of Gorenstein $(\mathcal{P}(R),\langle\mathcal{F}(R)\rangle)$-projective $R$-modules is special precovering. We state this as follows.
@@
Consider the definable duality pair $(\langle\mathcal{F}(R)\rangle,\langle\mathcal{I}(R^{\mathrm{op}})\rangle)$ in $\mathrm{Mod}(R)$, then the class $\mathcal{GP}_{\langle\mathcal{F}(R)\rangle}$ is the left part of a complete and hereditary cotorsion pair.
For each $n\geq 2$ there is a duality pair $(\mathcal{FP}_{n}\mbox{-}Flat(R),\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}}))$, which consists of the $FP_{n}$-flat left R-modules and $FP_{n}$-injective right $R$-modules [8, Theorems 5.5 & 5.6]. The right part $\mathcal{FP}_{n}\mbox{-}Inj(R^{\mathrm{op}})$ is definable [18, Example 2.21. (3)]. The classes of relative Gorenstein $R$-modules that comes from these duality pairs have been studied in [28] by A. Iacob, proving in a manner similar to that used in Proposition 3.6 that $\mathcal{GP}_{\mathcal{FP}_{n}\mbox{-}Flat(R)}$ is the left part of a complete and hereditary cotorsion pair, and thus is a special precovering class. From here, we declare the following result.